# Understanding Oversmoothing in Diffusion-Based GNNs From the Perspective of Operator Semigroup Theory

### v1

@@ 6 Experimental Verification
This section presents empirical experiments designed to validate the theoretical analysis of the oversmoothing problem in graph diffusion. Our objectives are twofold: (1) Verification of oversmoothing: We empirically demonstrate the oversmoothing phenomenon in certain graph diffusion structures as theorized, alongside testing a proposed method aimed at mitigating this issue in scenarios both with and without training, and across datasets characterized by differing levels of homophily; (2) Comparative Performance Analysis: The proposed method is evaluated against standard baselines in node classification benchmarks to illustrate its effectiveness in improving performance.
6.1 Specification on Ergodicity Breaking Terms
Following Theorem 4.5, we propose the following ergodicity breaking terms $\mathcal{C}$ w.r.t. the elementary function based on the attention matrix $A$:
  • Negative exponential: $e^{-A}=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}A^{n}$;
  • Exponential: $e^{A}=\sum_{n=0}^{\infty}\frac{1}{n!}A^{n}$;
  • Logarithm: $\log{(I+A)}=\sum_{n=1}^{\infty}\frac{(-1)^{n-1}}{n}A^{n}$;
  • Sine: $\sin{A}=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{(2n+1)!}A^{2n+1}$;
  • Cosine: $\cos{A}=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{(2n)!}A^{2n}$.
@@ 6.1 Specification on Ergodicity Breaking Terms
In the implementation level, we truncate the series by $N\in\mathbb{N}^{+}$, resulting breaking terms with different orders.
@@
Table 1: Node-classification results. Top three models are coloured by First, Second, Third
@@
CGNN $71.35\pm 4.05$ $74.31\pm 7.26$ $66.22\pm 7.69$ $35.95\pm 0.86$ $29.24\pm 1.09$ $46.89\pm 1.66$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}76.91\pm 1.81}$ $87.70\pm 0.49$ $87.10\pm 1.35$
@@
Sheaf ${85.40\pm 6.07}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}89.41\pm 4.74}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}84.86\pm 4.71}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}37.81\pm 1.15}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}56.34\pm 1.32}$ $67.71\pm 1.22$ $76.70\pm 1.57$ $89.49\pm 0.40$ $86.90\pm 1.13$
GRAFF $84.32\pm 5.51$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}85.68\pm 3.40}$ $71.08\pm 3.64$ $35.55\pm 1.20$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}57.60\pm 1.42}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}68.75\pm 1.83}$ ${75.27\pm 1.45}$ ${89.60\pm 0.40}$ ${87.66\pm 0.97}$
GREAD ${84.59\pm 4.53}$ $85.29\pm 4.49$ $73.78\pm 4.53$ $37.59\pm 1.06$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}58.62\pm 1.08}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}70.00\pm 1.70}$ $77.09\pm 1.78$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}90.01\pm 0.42}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}88.16\pm 0.81}$
Ours-Expn ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}88.65\pm 2.64}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}88.82\pm 3.62}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}76.48\pm 2.71}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}37.79\pm 0.86}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}50.83\pm 1.69}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}69.74\pm 1.26}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}77.48\pm 1.26}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}90.08\pm 0.49}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}88.31\pm 0.85}$
Ours-Exp ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}84.06\pm 4.59}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}87.64\pm 3.51}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}77.57\pm 3.83}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}37.37\pm 1.17}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}50.03\pm 1.90}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}69.41\pm 1.15}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}76.75\pm 1.76}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}89.92\pm 0.37}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}87.97\pm 1.09}$
Ours-Log ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}87.03\pm 5.24}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}86.67\pm 2.60}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}78.11\pm 3.90}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}37.55\pm 0.80}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}54.75\pm 1.57}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}70.75\pm 1.39}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}77.07\pm 1.60}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}90.02\pm 0.38}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}88.07\pm 1.24}$
Ours-Sin ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}85.41\pm 5.94}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}87.26\pm 5.21}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}77.84\pm 3.15}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}37.64\pm 0.89}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}56.35\pm 1.89}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}70.88\pm 1.58}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}77.13\pm 0.97}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}89.98\pm 0.45}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}88.21\pm 0.94}$
Ours-Cos ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}86.21\pm 5.20}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}85.69\pm 4.21}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}77.57\pm 1.73}$ ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}37.75\pm 0.89}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}52.44\pm 0.81}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}68.18\pm 1.69}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}77.21\pm 1.81}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}89.94\pm 0.39}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}88.33\pm 0.61}$
6.2 Oversmoothing Validation
@@
We first empirically validate the oversmoothing phenomenon as posited in our theoretical analysis. Dirichlet energy is utilized as the primary metric for this purpose, defining as:
$E\left(H,A\right)=\frac{1}{N}\sum_{u\in\mathcal{V}}\sum_{v\in\mathcal{N}(u)}a(u,v)\left\|h_{u}-h_{v}\right\|^{2}.$
@@
We adopt the synthetic Cora (Zhu et al., 2020) dataset for our experiments. To contextualize our findings, we compare the following models: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017), GRAND (Chamberlain et al., 2021b), GREAD (Choi et al., 2023) and ergodicity-breaking terms in Sec. 6.1, denoting as Ours-Exp, Ours-ExpN, Ours-Log, Ours-Sin, Ours-Cos. Detailed configurations and further results are in Appendix C.1.
@@
Question: Does the training process influence oversmoothing effects?
Answer: Figure 0(a) shows that the training process does not significantly impact the oversmoothing effects in various graph neural network models. Both trained and untrained versions of models such as GRAND-l, GRAND-nl, and the graph diffusion model with ergodicity-breaking terms (Ours-Exp, Ours-ExpN, Ours-Log, Ours-Sin, Ours-Cos) demonstrate similar trends in Dirichlet energy as the network depth increases. This consistency across models, regardless of their training status, suggests that oversmoothing is predominantly a function of the model’s inherent architecture rather than its training.
@@
Answer: Figure 0(b)’s error bands, which illustrate the variance due to different homophily levels, show that some models are more sensitive to homophily changes than others. For instance, models like GAT and GCN, with wider error bands, indicate a higher sensitivity to homophily levels, implying that their performance in managing oversmoothing varies more significantly with changes in homophily. Conversely, certain models from the ‘Ours’ series, characterized by narrower error bands, demonstrate greater stability against oversmoothing across varying homophily levels. This highlights that while some GNN models are robust against homophily variations, others may require careful consideration of the dataset’s homophily characteristics for optimal performance.
6.3 Node Classification

### v2

@@ 7. Experimental Verification
This section presents empirical experiments designed to validate the theoretical analysis of the oversmoothing problem in graph diffusion. Our objectives are twofold: (1) Verification of oversmoothing: We empirically demonstrate the oversmoothing phenomenon in certain graph diffusion structures as theorized, alongside testing the proposed method aimed at mitigating this issue in scenarios characterized by differing levels of homophily and test performance; (2) Comparative Performance Analysis: The proposed method is evaluated against standard baselines in node classification benchmarks to illustrate its effectiveness in improving performance.
7.1. Specification on Ergodicity Breaking Terms
@@
Following Theorem 7, we propose the following ergodicity breaking terms $\mathcal{C}$ w.r.t. the elementary function based on the adjacency matrix $A$: (1) Exponential: $e^{A}=\sum_{n=0}^{\infty}\frac{1}{n!}A^{n}$; (2) Negative exponential: $e^{-A}=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}A^{n}$; (3) Logarithm: $\log{(I+A)}=\sum_{n=1}^{\infty}\frac{(-1)^{n-1}}{n}A^{n}$; In the implementation level, we truncate the series by $N\in\mathbb{N}^{+}$, resulting breaking terms with different orders.
@@
Table 1. Node-classification results. Top three models are coloured by First, Second, Third
@@
CGNN $71.35\pm 4.05$ $74.31\pm 7.26$ $66.22\pm 7.69$ $35.95\pm 0.86$ $29.24\pm 1.09$ $46.89\pm 1.66$ ${\color[rgb]{0,0,0}76.91\pm 1.81}$ $87.70\pm 0.49$ $87.10\pm 1.35$
@@
GRAFF $84.32\pm 5.51$ ${\color[rgb]{0,0,0}85.68\pm 3.40}$ $71.08\pm 3.64$ $35.55\pm 1.20$ ${\color[rgb]{0,0,1}57.60\pm 1.42}$ ${\color[rgb]{0,0,0}68.75\pm 1.83}$ ${75.27\pm 1.45}$ ${89.60\pm 0.40}$ ${87.66\pm 0.97}$
GREAD ${\color[rgb]{1,.5,0}84.59\pm 4.53}$ $85.29\pm 4.49$ ${73.78\pm 4.53}$ ${\color[rgb]{0,0,1}37.59\pm 1.06}$ ${\color[rgb]{1,0,0}58.62\pm 1.08}$ ${\color[rgb]{0,0,1}70.00\pm 1.70}$ ${\color[rgb]{0,0,1}77.09\pm 1.78}$ ${\color[rgb]{1,.5,0}90.01\pm 0.42}$ ${\color[rgb]{0,0,1}88.16\pm 0.81}$
Ours-Expn ${\color[rgb]{1,0,0}88.65\pm 2.64}$ ${\color[rgb]{1,0,0}88.82\pm 3.62}$ ${\color[rgb]{1,.5,0}76.48\pm 2.71}$ ${\color[rgb]{1,0,0}37.79\pm 0.86}$ ${\color[rgb]{0,0,0}50.83\pm 1.69}$ ${\color[rgb]{1,.5,0}69.74\pm 1.26}$ ${\color[rgb]{1,0,0}77.48\pm 1.26}$ ${\color[rgb]{1,0,0}90.08\pm 0.49}$ ${\color[rgb]{1,0,0}88.31\pm 0.85}$
Ours-Exp ${\color[rgb]{0,0,0}84.06\pm 4.59}$ ${\color[rgb]{0,0,1}87.64\pm 3.51}$ ${\color[rgb]{0,0,1}77.57\pm 3.83}$ ${\color[rgb]{0,0,0}37.37\pm 1.17}$ ${\color[rgb]{0,0,0}50.03\pm 1.90}$ ${\color[rgb]{0,0,0}69.41\pm 1.15}$ ${\color[rgb]{0,0,0}76.75\pm 1.76}$ ${\color[rgb]{0,0,0}89.92\pm 0.37}$ ${\color[rgb]{0,0,0}87.97\pm 1.09}$
Ours-Log ${\color[rgb]{0,0,1}87.03\pm 5.24}$ ${\color[rgb]{1,.5,0}86.67\pm 2.60}$ ${\color[rgb]{1,0,0}78.11\pm 3.90}$ ${\color[rgb]{1,.5,0}37.55\pm 0.80}$ ${\color[rgb]{1,.5,0}54.75\pm 1.57}$ ${\color[rgb]{1,0,0}70.75\pm 1.39}$ ${\color[rgb]{1,.5,0}77.07\pm 1.60}$ ${\color[rgb]{0,0,1}90.02\pm 0.38}$ ${\color[rgb]{1,.5,0}88.07\pm 1.24}$
Table 2. Further node-classification results. The best results are highlighted in bold.
Computers Photo CoauthorCS ogbn-arxiv
GRAND $69.12\pm 0.41$ $84.64\pm 0.25$ $92.89\pm 0.30$ $91.27\pm 0.17$
GRAND++ $68.66\pm 0.63$ $84.99\pm 0.52$ $92.89\pm 0.23$ $91.51\pm 0.23$
ACMP $69.87\pm 0.33$ $85.30\pm 0.66$ $92.97\pm 0.41$ $91.23\pm 0.36$
GREAD $69.89\pm 0.45$ $85.81\pm 0.63$ $92.93\pm 0.13$ $90.93\pm 0.31$
Ours-Exp $70.35\pm 0.24$ $85.49\pm 0.98$ $93.71\pm 0.23$ $91.86\pm 0.20$
Ours-Expn $\mathbf{70.85\pm 0.24}$ $\mathbf{86.16\pm 0.54}$ $\mathbf{94.03\pm 0.31}$ $91.59\pm 0.32$
Ours-Log $70.68\pm 0.69$ $86.02\pm 0.31$ $93.75\pm 0.18$ $\mathbf{92.04\pm 0.30}$
7.2. Oversmoothing Validation
@@
We first empirically validate the oversmoothing phenomenon as posited in our theoretical analysis. Dirichlet energy (Maskey et al., 2023) is utilized as the primary metric for this purpose, defining as: $E\left(H,A\right)=\frac{1}{N}\sum_{u\in\mathcal{V}}\sum_{v\in\mathcal{N}(u)}a(u,v)\left\|\frac{h_{u}}{\sqrt{d_{u}}}-\frac{h_{v}}{\sqrt{d_{v}}}\right\|^{2},$ where $d_{u}$ is the degree of node $u$. We adopt the synthetic Cora (Zhu et al., 2020) dataset for our experiments.
@@
To contextualize our findings, we compare the following models: GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), GRAND (Chamberlain et al., 2021b), GREAD (Choi et al., 2023) and ergodicity-breaking terms in Sec. 7.1, denoting as Ours-Exp, Ours-ExpN, and Ours-Log. Detailed configurations are in Appendix C.1.
@@
Answer: Figure 2(b) illustrates the evolution of Dirichlet energy as the number of layers increases across different levels of homophily on synthetic Cora dataset. The results indicate that certain models exhibit greater sensitivity to changes in homophily than others. For instance, models such as GAT and GCN demonstrate a pronounced responsiveness to variations in homophily levels, suggesting that their performance in mitigating oversmoothing is considerably affected by these changes. In contrast, specific models within the ”Ours” series, characterized by closely aligned lines, exhibit enhanced stability against oversmoothing across varying levels of homophily. This observation underscores the notion that while some GNN models maintain robustness against fluctuations in homophily, others may necessitate careful consideration of the homophily characteristics inherent in the dataset to achieve optimal performance.
Question: How does the performance change with increasing network depth?
Answer: Figure 2(c) displays the test accuracy of different GNN models as a function of the number of layers on Cora dataset. The performance trends indicate a distinct variance among the models. For GCN and GAT, there is a marked decrease in performance as the number of layers increases, indicating a susceptibility to performance degradation due to oversmoothing. This effect is particularly pronounced beyond 32 layers where the test accuracy sharply declines. In contrast, for graph diffusion based models, performance remains relatively stable and high across all layer depths, suggesting these models are resilient to the deep layer effects, including potential oversmoothing issues. What’s more, the ”Ours” series methods achieve better performance than GRAND across the increasing layers, highlighting their effectiveness in handling deeper architectures.
7.3. Node Classification